Use elimination to solve:
3x-y=6
x-2y=-2

Answer by checkley71(8403)

3X-Y=6 X-2Y=-2 MULTIPLY THE SECOND Equation by -3 yields -3x+6y=6
-3x+6y=6 adding these two equations we get
---------
5y=12 or y=12/5
substituting this in the first equation we get
3x-12/5=6 or 3x=6+12/5 or 3x=42/5 or x=42/5times1/3 or x=14/5
proof -3(14/5)+6(12/5)=6 or -42/5+72/5=6 or (72-42)/5=6 or 30/5=6 or 6=6
and 3(14/5)-12/5=6 or 42/5-12/5=6 or (42-12)/5=6 or 30/5=6 or 6=6